OpenStudy (anonymous):

How to simplify log a^x + log a^9 + log a(x^2+1) - log a^5 ?

OpenStudy (anonymous):

\[\log_{a ^{x}} + \log_{a ^{9}} +\log_{a ^({x ^{2}+1})} -\log_{a ^{5}} \]

OpenStudy (gorv):

log has base a ?????

OpenStudy (gorv):

@jayjayokocha

OpenStudy (gorv):

log a^9- log a^5= \[\log_{a}9/5 \]

OpenStudy (gorv):

log a^x + log a(x^2+1)= \[\log_{a}((x^2+1)*(x)) \]

OpenStudy (gorv):

\[\log_{a}9/5 \] + \[\log_{a}(x^3+x) \] =\[\log_{a} ((x^3+x)*9/5)\]

OpenStudy (***[isuru]***):

Is it \[\log _{a} x + \log_{a} 9+ \log_{a} (x^{2}+1)+ \log_{a}5\] or \[\log a ^{x} + \log a {9} + \log a^{x^{2} + 1 } + \log a^{5}\] or \[\log _{a^{x}} + \log_{a^{9}} + \log_{a^{x^{2} +1}} + \log_{a^{5}}\]????

OpenStudy (***[isuru]***):

@jayjayokocha ??

OpenStudy (anonymous):

@***[ISURU]*** its the first equation. sorry for the confusion